# Wasserstein Spatial Depth

### v1

@@ A.2 Transformation invariance
Theorem 1.2 in [23] describes the group of isometries of $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ for $d>2$. Any isometry $F$, can be written as the composition of $\Phi(\varphi)$ and a trivial isometry. Recall that $\Phi(\varphi):P\mapsto\Phi(\varphi)(P)$ where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a linear isometry and $\Phi(\varphi)(P)$ is the law of the random variable
@@
$S_{P}(\varphi(\mathbf{x}-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}])+\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]))=\varphi(T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])$
@@
$=1-\bigg{(}\int\bigg{\|}\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\varphi(\mathbf{x}-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}])+\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]))}{\mathcal{W}_{2}(P,Q)}$
@@
$1-\bigg{(}\int\bigg{\|}\varphi\left(\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])}{\mathcal{W}_{2}(P,Q)}\bigg{]}\right)$
@@ A.3 Vanishing at infinity
Note that $\mathcal{W}_{2}(Q_{n},P)\to\infty$ implies that for any other $P^{\prime}\in\mathcal{P}_{2}(P)$,
@@
with the convention $\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}=0$ if ${\mathcal{W}_{2}(P,Q_{n})}=0$. First we want to get rid of this last pathological case. Let
$A_{n}:=\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x}).$ (11)
Let $E_{n}=\{Q_{n}\}$. Note that when $P\in E_{n}$, a 00 appears in the expression of $A_{n}$ (recall the convention $/0=0$ 0). For each $n$, we modify ${\bf P}={\bf P}_{1}+{\bf P}_{2}$, where ${\bf P}_{1}$ is a measure on $\mathcal{P}_{2}(\mathbb{R}^{d})\backslash E_{n}$ and ${\bf P}_{2}$ is a measure on $E_{n}$, by ${\bf P}^{\prime}={\bf P}_{1}+{\bf\widetilde{P}}_{2}$, where ${\bf\widetilde{P}}_{2}$ is an arbitrary measure on $\mathcal{P}_{2}(\mathbb{R}^{d})\backslash E_{n}$ such that ${\bf\widetilde{P}}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))={\bf P}(E_{n})$. Note that ${\bf P}^{\prime}$ is also a probability measure.
@@
We prove that $A_{n}\to 1$, where $A_{n}$ is defined in (11). To do so, let $P^{\prime}$ be an independent copy of $P$, so that
@@
$=\int{\mathbb{E}_{P,P^{\prime}\sim\mathbf{P}}\left[\frac{\left\langle\mathbf{x}-T_{Q_{n},P}(\mathbf{x}),\mathbf{x}-T_{Q_{n},P^{\prime}}(\mathbf{x})\right\rangle}{\mathcal{W}_{2}(P,Q_{n})\mathcal{W}_{2}(P^{\prime},Q_{n})}\right]}\mathrm{d}Q_{n}(\mathbf{x}).$ (12)
@@
$C_{n}(P,P^{\prime}):=\frac{\left\langle B_{P,n},B_{P^{\prime},n}\right\rangle_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\leq\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}=1,$
@@
$C_{n}(P,P^{\prime})\longrightarrow 1,\quad\text{for }{\bf P}\otimes{\bf P}-{\rm a.e.}\ (P,P^{\prime}).$ (13)
@@
$\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}}\frac{\|T_{Q_{n},P}\|_{L^{2}(Q_{n})}+\|T_{Q_{n},P^{\prime}}\|_{L^{2}(Q_{n})}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\\ +\left|\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},B_{P,n}\left(\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})}\right|.$ (14)
The first term of (14) tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. Indeed, using Cauchy-Schwarz inequality, $(a+b)^{2}\leq 2(a^{2}+b^{2})$ and the equality $\|T_{Q_{n},P}\|_{L^{2}(Q_{n})}^{2}=\int\|\mathbf{x}\|^{2}\mathrm{d}P(\mathbf{x})$, the following upper bound
$\frac{\int\|\mathbf{x}\|^{2}\mathrm{d}P(\mathbf{x})+\int\|\mathbf{x}\|^{2}\mathrm{d}P^{\prime}(\mathbf{x})}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}$ (15)
holds for the first term of (14). The latter clearly tends to 00 since $\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}=\mathcal{W}_{2}(Q_{n},P^{\prime})$.
To show that the second term of (14) also tends to 00 we use the bound
$\left|\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},B_{P,n}\left(\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})}\right|\\ \leq\left|\|B_{P,n}\|\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}|}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right|$
@@
$\left|\|B_{P,n}\|\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}|}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right|$ $=\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}|}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)$
@@
The latter can be upper bounded by (15), so that the second term of (14) also tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. This implies $C_{n,2}(P,P^{\prime})$ tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. Hence, (13) holds.
@@
$\mathcal{W}_{2}^{2}(P,Q)=\inf_{\pi\in\Pi(P,Q)}\frac{1}{2}\int\|\mathbf{x}-\mathbf{y}\|^{2}\mathrm{d}\pi(\mathbf{x},\mathbf{y})$ (16)
@@
$\mathcal{W}_{2}^{2}(P,Q)=\sup_{(f,g)\in\Phi}\left\{\int f(\mathbf{x})\,\mathrm{d}Q(\mathbf{x})+\int g(\mathbf{y})\,\mathrm{d}P(\mathbf{y})\right\},$ (17)
where $\Phi=\{(f,g)\in\mathcal{C}(\mathbb{R}^{d})\times\mathcal{C}(\mathbb{R}^{d}):\ f(\mathbf{x})+g(\mathbf{y})\leq\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|^{2}\}$. Here $\mathcal{C}(\mathbb{R}^{d})$ is the set of continuous functions on $\mathbb{R}^{d}$. We denote as $(f_{Q,P},f_{P,Q})$ the solutions of (17). It is well-known that $\nabla f_{Q,P}(\mathbf{x})=\mathbf{x}-T_{Q,P}(\mathbf{x})$. The indicator function of a set $A$ is denoted as ${\bf 1}_{A}$. Now we argue by contradiction. We assume first that there exists
$Q\in\mathcal{P}_{2}^{a.c}(\mathbb{R}^{d})\cap\operatorname*{arg\,min}_{P}\mathbb{E}_{P\sim\mathbf{P}}[\mathcal{W}_{2}(P,Q)]$
with $Q\not\in\{P_{1},\ldots,P_{n}\}$ and we assume that the set $\mathcal{K}$ of all $\mathbf{x}$ such that
@@
has positive measure $Q(\mathcal{K})>0$. As $T_{Q,P}$ is the gradient of a l.s.c. convex function, it is continuous $Q$-a.e., so that $s$ is also continuous $Q$-a.e. Therefore, there exists a compact convex set with non-empty interior $U$ such that $U\subset\mathcal{K}$. Consider the signed measure $h$ such that $\frac{dh}{dQ}=-{\bf 1}_{U}\left(f_{Q,P}-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)$. Note that $h(\mathbb{R}^{d})=0$ and $Q+th$ is a probability measure with finite second order moment for all $t$ in a neighborhood of zero. Since $(\cdot)^{1/2}$ is concave,
@@
Using the dual formulation (17) we obtain for $t$ in a neighborhood of zero,
@@
$\frac{\mathcal{W}_{2}(P,Q+th)-\mathcal{W}_{2}(P,Q)}{t}\\ \leq-\frac{\int_{U}(f_{Q,P}(\mathbf{x})-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))(f_{Q+th,P}(\mathbf{x})-\int_{U}f_{Q+th,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))\mathrm{d}Q(\mathbf{x})}{2\mathcal{W}_{2}(P,Q)}.$
@@
$M(P):=\frac{1}{2}\frac{\int_{U}(f_{Q,P}(\mathbf{x})-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))^{2}\mathrm{d}Q(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}$
@@
$\|\phi\|_{U}:=\left(\int_{U}\left(\phi(\mathbf{x})-\int_{U}\phi(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)^{2}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
@@
Since $s(\mathbf{x})\neq 0$ for $\mathbf{x}\in U$, the function $U\ni\mathbf{x}\mapsto\mathbb{E}[f_{Q,P}(\mathbf{x})]$ is non constant, which implies that
$\mathbb{E}_{P\sim{\bf P}}[M(P)]:=\frac{1}{2}\mathbb{E}_{P\sim{\bf P}}\left[\frac{\int_{U}(f_{Q,P}(\mathbf{x})-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))^{2}\mathrm{d}Q(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]>0.$
@@
$\mathbb{E}_{P\sim{\bf P}}\left[\frac{\|f_{Q,P}\|_{U}\|f_{Q,P}-f_{Q+th,P}\|_{U}}{\mathcal{W}_{2}(P,Q)}\right]\to 0\quad{\rm as}\ t\to 0,$ (18)
@@
As $\mathcal{W}_{2}(Q_{n},Q)\to 0$ and $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is atomless there exists an open Wasserstein ball
@@
${\bf P}(\mathcal{V}_{\beta}^{c})\leq\epsilon.$ (19)
Moreover, we can assume that $n$ is large enough such that $\mathcal{W}_{2}(Q_{n},Q)\leq\beta/2$, which implies that
$\mathcal{W}_{2}(Q_{n},P)\geq\mathcal{W}_{2}(P,Q)-\mathcal{W}_{2}(Q_{n},Q)\geq\beta/2,$ (20)
@@
so that, arguing as in Subsection A.1 and using (19), we derive the bound $|A_{n}-B_{n}|\leq\epsilon$ for all $n\in\mathbb{N}$. By the same means $|A-B|\leq\epsilon$ where
@@
and thus (21) holds. The continuous mapping theorem with the function $(\mathbf{x},\mathbf{y},\mathbf{z})\mapsto(\mathbf{y}-\mathbf{x},\mathbf{z}-\mathbf{x})$ implies that
@@
for all $P,P^{\prime}$ such that $\mathcal{W}_{2}(Q,P)>0$ and $\mathcal{W}_{2}(Q,P^{\prime})>0$. As a consequence, (22) holds for ${\bf P}$-a.e. $P,P^{\prime}$.
@@
where the above limit holds due to dominated convergence
@@
we only need to prove that $\mathbf{Y}_{n}$ is uniformly integrable. The bound (20) implies that it is enough to show that each of the terms of the right hand side of
$|\left\langle\mathbf{X}_{n}-T_{Q_{n},P}(\mathbf{X}_{n}),\mathbf{X}_{n}-T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\right\rangle|\\ \leq\|\mathbf{X}_{n}\|^{2}+\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|+\|T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|+\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\|$ (23)
@@
$\lim_{R\to+\infty}\sup_{P\in\mathcal{V}_{\beta}}\int_{{\|\mathbf{x}\|^{2}>R}}\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})=0$ (24)
@@
$\lim_{R\to+\infty}\sup_{n\in\mathbb{N}}\int_{{\|\mathbf{x}\|^{2}>R}}\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})=0.$ (25)
The last limit (25) implies that the sequence $\{\|\mathbf{X}_{n}\|^{2}\}_{n\in\mathbb{N}}$ is uniformly integrable, so that the first term of the right-hand-side of (23) is uniformly integrable. For the second, we observe that
@@
$\qquad\qquad+\left(\mathbb{E}\left[\|T_{Q_{n},P}(\mathbf{X}_{n})\|{\bf 1}_{\|T_{Q_{n},P}(\mathbf{X}_{n})\|>R^{\frac{1}{2}}}\right]\mathbb{E}\left[\|\mathbf{X}_{n}\|^{2}\right]\right)^{\frac{1}{2}}$

### v2

@@ A.2 Transformation invariance
Theorem 1.2 in [32] describes the group of isometries of $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ for $d\geq 2$. Any isometry $F$ can be written as the composition of $\Phi(\varphi)$ and a trivial isometry. Recall that $\Phi(\varphi):P\mapsto\Phi(\varphi)(P)$ where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a linear isometry and $\Phi(\varphi)(P)$ is the law of the random variable
@@
$S_{P}(\varphi(\mathbf{x}-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}])+\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}])=\varphi(T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]$
@@
$=1-\bigg{(}\int\bigg{\|}\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\varphi(\mathbf{x}-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}])+\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]}{\mathcal{W}_{2}(P,Q)}$
@@
$=$ $1-\bigg{(}\int\bigg{\|}\varphi\left(\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])}{\mathcal{W}_{2}(P,Q)}\bigg{]}\right)$
@@ A.3 Vanishing at infinity
Note that $\mathcal{W}_{2}(Q_{n},P)\to\infty$ implies that for any other $P^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@
with the convention $\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}=\boldsymbol{0}$ if ${\mathcal{W}_{2}(P,Q_{n})}=0$. First we want to get rid of this last pathological case. Let
$A_{n}:=\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x}).$ (A.1)
[…]
@@
We prove that $A_{n}\to 1$, where $A_{n}$ is defined in (A.1). To do so, let $P^{\prime}$ be an independent copy of $P$, so that
@@
$=\int{\mathbb{E}_{P,P^{\prime}\sim\mathbf{P}}\left[\frac{\left\langle\mathbf{x}-T_{Q_{n},P}(\mathbf{x}),\mathbf{x}-T_{Q_{n},P^{\prime}}(\mathbf{x})\right\rangle}{\mathcal{W}_{2}(P,Q_{n})\mathcal{W}_{2}(P^{\prime},Q_{n})}\right]}\mathrm{d}Q_{n}(\mathbf{x}).$ (A.2)
@@
$\left|C_{n}(P,P^{\prime})\right|:=\left|\frac{\left\langle B_{P,n},B_{P^{\prime},n}\right\rangle_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right|\leq\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}=1,$
@@
$C_{n}(P,P^{\prime})\longrightarrow 1,\quad\text{for }{\bf P}\otimes{\bf P}-{\rm a.e.}\ (P,P^{\prime}).$ (A.3)
@@
$\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}}\frac{\|T_{Q_{n},P}\|_{L^{2}(Q_{n})}+\|T_{Q_{n},P^{\prime}}\|_{L^{2}(Q_{n})}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\\ +\left|\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},B_{P,n}\left(\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})}\right|.$ (A.4)
The first term of (A.4) tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. Indeed, using the equality $\|T_{Q_{n},P}\|_{L^{2}(Q_{n})}^{2}=\int\|\mathbf{x}\|^{2}\mathrm{d}P(\mathbf{x})$, the first term of (A.4) is equal to
$\frac{\sqrt{\int\|\mathbf{x}\|^{2}\mathrm{d}P(\mathbf{x})}+\sqrt{\int\|\mathbf{x}\|^{2}\mathrm{d}P^{\prime}(\mathbf{x})}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}.$ (A.5)
The latter clearly tends to 00 since $\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}=\mathcal{W}_{2}(Q_{n},P^{\prime})$.
To show that the second term of (A.4) also tends to 00 we use the bound
$\left|\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},B_{P,n}\left(\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})}\right|\\ \leq\left|\|B_{P,n}\|\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right|$
@@
$\left|\|B_{P,n}\|\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right|$ $=\left|\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right|$
@@
The latter can be upper bounded by (A.5), so that the second term of (A.4) also tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. This implies $C_{n,2}(P,P^{\prime})$ tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. Hence, (A.3) holds.
@@
$\mathcal{W}_{2}^{2}(P,Q)=\inf_{\pi\in\Pi(P,Q)}\frac{1}{2}\int\|\mathbf{x}-\mathbf{y}\|^{2}\mathrm{d}\pi(\mathbf{x},\mathbf{y})$ (B.1)
@@
$\mathcal{W}_{2}^{2}(P,Q)=\sup_{(f,g)\in\Phi}\left\{\int f(\mathbf{x})\,\mathrm{d}Q(\mathbf{x})+\int g(\mathbf{y})\,\mathrm{d}P(\mathbf{y})\right\},$ (B.2)
where $\Phi=\{(f,g)\in\mathcal{C}(\mathbb{R}^{d})\times\mathcal{C}(\mathbb{R}^{d}):\ f(\mathbf{x})+g(\mathbf{y})\leq\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|^{2}\}$. Here $\mathcal{C}(\mathbb{R}^{d})$ is the set of continuous functions on $\mathbb{R}^{d}$. We denote as $(f_{Q,P},f_{P,Q})$ the solutions of (B.2). It is well-known that $\nabla f_{Q,P}(\mathbf{x})=\mathbf{x}-T_{Q,P}(\mathbf{x})$. Now we argue by contradiction. We assume first that there exists
$Q\in\mathcal{P}_{2}^{a.c}(\mathbb{R}^{d})\cap\operatorname*{arg\,min}_{Q^{\prime}}\mathbb{E}_{P\sim\mathbf{P}}[\mathcal{W}_{2}(P,Q^{\prime})]$
with $Q\not\in\{P_{1},\ldots,P_{n}\}$ and we assume that the set $\mathcal{K}^{\prime}$ of all $\mathbf{x}$ such that
@@
has positive measure $Q(\mathcal{K}^{\prime})>0$. As $T_{Q,P}$ is the gradient of a lower semi continuous convex function, it is continuous $Q$-a.e., so that $s$ is also continuous $Q$-a.e. Therefore, there exists a compact convex set with non-empty interior $U$ such that $U\subset\mathcal{K}^{\prime}$. Consider the signed measure $h$ such that $\frac{dh}{dQ}=-{\bf 1}_{U}\left(f_{Q,P}-\frac{1}{Q(U)}\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)$, where ${\bf 1}_{U}$ is the indicator function of the set $U$.
Note that $h(\mathbb{R}^{d})=0$ and $Q+th$ is a probability measure with finite second order moment for all $t$ in a neighborhood of zero. Since $(\cdot)^{1/2}$ is concave,
@@
Using the dual formulation (B.2) we obtain for $t$ in a neighborhood of zero,
@@
$\frac{\mathcal{W}_{2}(P,Q+th)-\mathcal{W}_{2}(P,Q)}{t}\\ \leq-\frac{\int_{U}\left(f_{Q,P}(\mathbf{x})-\frac{1}{Q(U)}\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)\left(f_{Q+th,P}(\mathbf{x})-\frac{1}{Q(U)}\int_{U}f_{Q+th,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)\mathrm{d}Q(\mathbf{x})}{2\mathcal{W}_{2}(P,Q)}.$
@@
$M(P):=\frac{1}{2}\frac{\int_{U}\left(f_{Q,P}(\mathbf{x})-\frac{1}{Q(U)}\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)^{2}\mathrm{d}Q(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}$
@@
$\|\phi\|_{U}:=\left(\int_{U}\left(\phi(\mathbf{x})-\frac{1}{Q(U)}\int_{U}\phi(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)^{2}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
@@
Since $s(\mathbf{x})\neq 0$ for $\mathbf{x}\in U$, the function $U\ni\mathbf{x}\mapsto\mathbb{E}_{P\sim{\bf P}}[f_{Q,P}(\mathbf{x})]$ is non constant, which implies that
$\mathbb{E}_{P\sim{\bf P}}[M(P)]:=\frac{1}{2}\mathbb{E}_{P\sim{\bf P}}\left[\frac{\int_{U}\left(f_{Q,P}(\mathbf{x})-\frac{1}{Q(U)}\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)^{2}\mathrm{d}Q(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]>0.$
@@
$\mathbb{E}_{P\sim{\bf P}}\left[\frac{\|f_{Q,P}\|_{U}\|f_{Q,P}-f_{Q+th,P}\|_{U}}{\mathcal{W}_{2}(P,Q)}\right]\to 0\quad{\rm as}\ t\to 0,$ (B.3)
@@
As $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is atomless there exists an open Wasserstein ball
@@
${\bf P}(\mathcal{V}_{\beta}^{c})\leq\epsilon.$ (C.1)
Moreover, as $\mathcal{W}_{2}(Q_{n},Q)\to 0$, we can assume that $n$ is large enough such that $\mathcal{W}_{2}(Q_{n},Q)\leq\beta/2$, which implies that
$\mathcal{W}_{2}(Q_{n},P)\geq\mathcal{W}_{2}(P,Q)-\mathcal{W}_{2}(Q_{n},Q)\geq\beta/2,$ (C.2)
@@
so that, arguing as in Subsection A.1 and using (C.1), we derive the bound $|A_{n}-B_{n}|\leq\epsilon$ for all $n\in\mathbb{N}$. By the same means $|A-B|\leq\epsilon$ where
@@
and thus (C.3) holds. The continuous mapping theorem with the function $(\mathbf{x},\mathbf{y},\mathbf{z})\mapsto(\mathbf{y}-\mathbf{x},\mathbf{z}-\mathbf{x})$ implies that
@@
for all $P,P^{\prime}$ such that $\mathcal{W}_{2}(Q,P)>0$ and $\mathcal{W}_{2}(Q,P^{\prime})>0$. As a consequence, (C.4) holds for ${\bf P}$-a.e. $P,P^{\prime}$.
@@
where the above limit holds due to dominated convergence.
@@
we only need to prove that $\mathbf{Y}_{n}$ is uniformly integrable. The bound (C.2) implies that it is enough to show that each of the terms of the right hand side of
$|\left\langle\mathbf{X}_{n}-T_{Q_{n},P}(\mathbf{X}_{n}),\mathbf{X}_{n}-T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\right\rangle|\\ \leq\|\mathbf{X}_{n}\|^{2}+\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|+\|T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|+\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\|$ (C.5)
@@
$\lim_{R\to+\infty}\sup_{P\in\mathcal{V}_{\beta}}\int_{{\|\mathbf{x}\|^{2}>R}}\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})=0$ (C.6)
@@
$\lim_{R\to+\infty}\sup_{n\in\mathbb{N}}\int_{{\|\mathbf{x}\|^{2}>R}}\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})=0.$ (C.7)
The last limit (C.7) implies that the sequence $\{\|\mathbf{X}_{n}\|^{2}\}_{n\in\mathbb{N}}$ is uniformly integrable, so that the first term of the right-hand-side of (C.5) is uniformly integrable. For the second, we observe that
@@
$\qquad\qquad+\left(\mathbb{E}\left[\|T_{Q_{n},P}(\mathbf{X}_{n})\|^{2}{\bf 1}_{\|T_{Q_{n},P}(\mathbf{X}_{n})\|>R^{\frac{1}{2}}}\right]\mathbb{E}\left[\|\mathbf{X}_{n}\|^{2}\right]\right)^{\frac{1}{2}}$